### v1

# DLTPose: 6DoF Pose Estimation From Accurate Dense Surface Point Estimates

Abstract
We propose DLTPose, a novel method for 6DoF object pose estimation from RGBD images that combines the accuracy of sparse keypoint methods with the robustness of dense pixel-wise predictions. DLTPose predicts per-pixel radial distances to a set of minimally four keypoints, which are then fed into our novel Direct Linear Transform (DLT) formulation to produce accurate 3D object frame surface estimates, leading to better 6DoF pose estimation. Additionally, we introduce a novel symmetry-aware keypoint ordering approach, designed to handle object symmetries that otherwise cause inconsistencies in keypoint assignments. Previous keypoint-based methods relied on fixed keypoint orderings, which failed to account for the multiple valid configurations exhibited by symmetric objects, which our ordering approach exploits to enhance the model’s ability to learn stable keypoint representations. Extensive experiments on the benchmark LINEMOD, Occlusion LINEMOD and YCB-Video datasets show that DLTPose outperforms existing methods, especially for symmetric and occluded objects, demonstrating superior Mean Average Recall values of 86.5% (LM), 79.7% (LM-O) and 89.5% (YCB-V). The code is available at https://anonymous.4open.science/r/DLTPose_/.
1 Introduction
Object pose estimation is a fundamental problem in computer vision with broad applications in robotics, augmented reality, and autonomous systems [Peng, Tremblay, Pavlakos]. The goal is to determine an object’s six-degree-of-freedom (6DoF) pose, which comprises both a 3D rotation and a 3D translation, from visual data. This task is particularly challenging due to factors such as occlusions, background clutter, sensor noise, varying lighting conditions, and object symmetries, all of which introduce ambiguities and uncertainties.
Figure 1: Visualization of DLTPose surface estimation. (a) For image point $p_{i}$, four radial distances $\hat{r}_{1},\hat{r}_{2},\hat{r}_{3},\hat{r}_{4}$ are estimated per-pixel, as the Euclidean distance to four predefined keypoints $k_{j}$. The DLT solution uses these keypoints and radial distances to estimate object frame 3D surface points $\bar{p}_{i}$. (b) Estimated surface points (blue) overlaid on the object mesh, in the object frame.
Among modern approaches, two dominant paradigms have emerged: sparse and dense methods [s24041076]. Sparse methods  [Wu, He, Peng] focus on predicting a small set of keypoints with high accuracy, which ultimately leads to increased accuracy in estimated poses. While they can be highly accurate, one limitation of sparse methods is that their reliance on a limited number of keypoints makes them more susceptible to occlusions, as missing or misidentified keypoints can significantly impact overall estimation. In contrast, dense methods  [Brachmann, Park, Haugaard, labbe2020] predict per-pixel 3D surface coordinates in the object frame using 2D image inputs, providing high robustness through redundant predictions. This redundancy helps mitigate errors by utilizing techniques such as RANSAC to filter out inconsistent and inaccurate predictions. While the increased number of surface points estimated can increase accuracy in occluded and cluttered scenes, dense methods do not specifically focus on obtaining highly accurate estimates of the 3D visible surface of the object. Instead, their primary objective is to generate a broad and redundant representation, often at the expense of accurately localizing individual points.
This work proposes a unified approach that combines the accuracy of sparse with the redundancy of dense methods. Our method, called DLTPose, trains a CNN on RGB-D data to estimate a per-pixel (minimally) four-dimensional representation, where each channel encodes a radial distance, defined as the Euclidean distance between a 3D scene point corresponding to a 2D image pixel, relative to a predefined 3D keypoint. A minimum of four keypoints is required to estimate the pixel surface coordinates in the object-centric reference frame, by solving a novel and highly accurate Direct Linear Transform (DLT) formulation. In this way, the radial quantities inferred from the network are combined with the DLT method to produce an accurate 2D-3D estimate of the visible surface of the object. The estimated object frame points, along with their corresponding 3D camera frame points, are then passed to a RANSAC-enabled Umeyama algorithm [umeyama] to estimate the final object pose. This pipeline enables precise and robust dense 3D surface estimation, ultimately leading to improved pose accuracy compared to previous methods where the dense 2D-3D surface points are inferred using end-to-end networks, albeit with lower accuracy [Park, labbe2020, Haugaard].
By integrating the precision of sparse keypoint-based methods with the redundancy of dense approaches, the proposed framework enhances both robustness and accuracy in object pose estimation. This unified strategy improves the fidelity of 3D surface reconstruction, which in turn leads to more reliable pose estimation, particularly in occluded scenarios where conventional methods struggle.
Handling objects with inherent symmetries is challenging, as visually identical orientations can mislead network training and prediction. Standard loss functions often fail to address these symmetries, resulting in large errors when equivalent poses are treated as distinct. Previous methods attempt to mitigate this by restricting training poses [kehl, Markus] or mapping predictions to the nearest symmetric equivalent [Mahdi], but these approaches struggle with discrete symmetries, leading to reduced generalization and inconsistencies. Instead of limiting training diversity or applying post-hoc corrections, our approach inherently incorporates symmetry awareness by ensuring keypoints remain stable across symmetrical transformations. Unlike traditional keypoint-based methods, which assign fixed keypoints and suffer from inconsistencies under symmetric rotations, our method maintains stable keypoint relationships, improving pose accuracy and robustness across different viewpoints.
The key contributions of this paper are as follows:
  • We propose a novel DLT formulation to estimate accurate 3D object frame surface points from 2D image points by leveraging per-pixel radial distance predictions to a minimal set of four keypoints. The resulting accurate 3D surface estimates improve the accuracy of pose estimation, especially in difficult cluttered and occluded scenarios.
  • We introduce a symmetry-aware keypoint framework that dynamically reorders keypoints based on their relative position in the camera view, ensuring stable radial map learning across symmetric transformations. This prevents inconsistencies in keypoint assignments, reducing regression errors and improving robustness in pose estimation for symmetric objects.
  • Our method achieves state-of-the-art results compared to recent leading methods on benchmark datasets, demonstrating superior performance in handling occlusions, cluttered environments, and symmetric objects.
2 Literature Review
Deep learning-based pose estimation techniques can be broadly classified into sparse and dense strategies [s24041076], each with distinct advantages and challenges. Sparse methods rely on detecting a small set of keypoints in an image and establishing correspondences with known 3D points defined in the object frame. BB8 [rad2017bb8] was an early learning-based sparse method that regressed 2D projections of objects’ 3D bounding box corners, and then applied a PnP solver for pose estimation. KeyPose [sundermeyer2020augmented] introduced a structured approach to keypoint regression, leveraging data augmentation techniques to improve robustness. It extended the idea of bounding box-based keypoint detection by incorporating keypoint refinement strategies, making it more robust to perspective distortions. PVNet [Peng] regressed vector fields pointing toward keypoint locations, relying on a RANSAC-based voting strategy to filter out noisy predictions. Building on PVNet, PVN3D [He] extended keypoint regression to 3D by voting on clusters of offsets to keypoints within point cloud data. RCVPose [Wu] extended the sparse approach by accumulating votes on the surface of intersecting spheres defined by radial values regressed for each image pixel. It was shown that the 1D radial value resulted in more accurately localized keypoints compared against the 2D vector value of PVNet and the 3D offset value of PVN3D. Despite these advances, most sparse methods relied on heuristically chosen keypoints that may not generalize well across different object shapes. KeyGNet [wu2024keygnet] addressed this limitation by learning optimal keypoint locations. However, sparse methods remain susceptible to failures when keypoints are missing or misidentified due to occlusions or viewpoint changes.
Dense methods addressed these limitations by estimating per-pixel 3D surface points of objects, providing richer geometric information for pose estimation. In the early seminal work of [Brachman1], a Random Forest was trained to estimate dense 3D surface coordinates in the object reference frame, which in later work was refined with an energy minimization approach using a DNN [Brachmann]. In the Normalized Object Coordinate Space (NOCS) framework [wang2019normalized], a canonical representation was proposed to estimate dense surface coordinates by normalizing object shape and scale, thereby improving learning generalization across different objects. Pix2Pose [Park] built upon this concept by regressing per-pixel object coordinate predictions through a fully convolutional backbone, refining the resulting RANSAC PnP derived pose estimates through the use of confidence maps.
Further building upon the dense approach, DenseFusion [wang2019densefusion] introduced a dense feature fusion mechanism that combined RGB and depth features at the per-pixel level, leveraging both color and geometric information and allowing for more robust 6DoF pose estimation in cluttered and occluded environments. Similarly, DPOD [zakharov2019dpod] formulated pose estimation as a dense object coordinate regression problem, where each pixel directly predicted its corresponding 3D model coordinate. By leveraging a pre-trained deep feature extractor, DPOD improved prediction consistency, particularly for objects with complex geometries. Coupled Iterative Refinement [lipson2022coupled] introduced a multi-stage process that refined pose predictions by progressively aligning estimated 3D coordinates with observed depth data, mitigating initial pose estimation errors and enhancing robustness in challenging scenes. SurfEmb [Haugaard] introduced a surface embedding-based approach that learned a continuous representation of object surfaces instead of discrete coordinate mappings. This method improved surface correspondence estimation by embedding geometric features in a high-dimensional space, allowing for more accurate object localization, even in occluded or textureless regions.
Hybrid approaches attempt to integrate the benefits of both sparse and dense methods by leveraging the accuracy of keypoints while incorporating dense feature representations. FFB6D [he2021ffb6d] fuses feature-based keypoint detection with dense depth representations to improve robustness in cluttered and occluded environments. Despite these improvements, hybrid methods still struggle with symmetric objects and accurate surface point localization. Our method leverages elements of both sparse and dense methods by integrating per-object pixel 3D surface estimates from a set of (minimally four) keypoints, using a novel Direct Linear Transform (DLT) formulation. Our framework is similar to that of Pix2Pose, with the increased accuracy of our DLT surface estimation and our treatment of object symmetries leading to improved pose estimation accuracy.
Figure 2: DLT network architecture. Both training and inference estimate radial map $\hat{R}$, which has channel depth $N_{k}\geq 4$ corresponding to the number of keypoints. At least 4 keypoints are required to solve the DLT formulation (Eq. 2) to estimate object surface points $\mathbf{\widehat{\overline{X}}}$.
3 Method
3.1 DLT Surface Estimation
In keypoint methods, quantities regressed for each foreground image pixel are aggregated to estimate a sparse set of keypoints in the camera frame. For example, in RCVPose [Wu], the radial distances from each pixel to each keypoint are inferred, and are combined with the pixels’ corresponding depth values to vote upon the keypoints’ 3D coordinates in the scene frame. In this work, we propose an inverse process, whereby the radial quantities inferred at each image pixel are combined with known object frame keypoint coordinates, to estimate the corresponding object frame 3D pixel coordinates. In this way, the same inferred radial values that were previously used to estimate image frame keypoints [Wu], are repurposed here for 3D surface reconstruction in the object frame.
[…]
At inference, a network estimates the radial (i.e. Euclidean) distance $\widehat{r}_{j}$ between each $p$ and $k_{j}$. As rigid transformations are isometric and preserve distances, this radial value is therefore also an estimate of the distance between corresponding object frame surface points and keypoints:
$\widehat{r}_{j}\simeq r_{j}=|\!|p-k_{j}|\!|=|\!|\overline{p}-\overline{k}_{j}|\!|.$ (1)
Expanding Eq. 1 and collecting terms (as derived in Sec. S.2 of the Supplementary Material) gives:
$\begin{bmatrix}\!\!-2\overline{x}_{k_{1}}\!\!\!\!\!\!\!\!&\!\!\!\!-2\overline{y}_{k_{1}}\!\!\!\!\!\!\!\!&\!\!\!\!\!\!\!\!-2\overline{z}_{k_{1}}\!\!\!\!\!\!\!\!&\!\!\!\!1&\!\!\!\!(|\!|\overline{k}_{1}|\!|^{2}\!-\!\widehat{r}_{1}^{2})\\ \vdots&\vdots&\vdots&\!\!\!\!\vdots&\vdots\\ \!-2\overline{x}_{k_{N_{k}}}\!\!\!\!\!\!\!\!&-2\overline{y}_{k_{N_{k}}}\!\!\!\!\!\!\!\!&-2\overline{z}_{k_{N_{k}}}\!\!\!\!&\!\!\!\!1&\!\!\!\!(|\!|\overline{k}_{N_{k}}|\!|^{2}\!-\!\widehat{r}_{N_{k}}^{2})\\ \end{bmatrix}\begin{bmatrix}\overline{x}\\ \overline{y}\\ \overline{z}\\ |\!|\overline{p}|\!|^{2}\\ 1\end{bmatrix}=0.$ (2)
Eq. 2 is in the familiar $\mathbf{A\overline{X}}\!\!=\!\!0$ form of the Direct Linear Transform (DLT) [Hartley:2003:MVG:861369]. Matrix $\mathbf{A}$ comprises the known coordinates of the object frame keypoints $\{\overline{k}_{j}\}_{j=1}^{N_{k}}$ and their corresponding radial values $\hat{r}_{j}$ inferred from image frame point $p$, whereas $\mathbf{\overline{X}}$ comprises the unknown object frame coordinates $\overline{p}$ of $p$. Composing $\mathbf{A}$ from $N_{k}\geq 4$ non-coplanar keypoints and their respective radial estimates and performing Singular Value Decomposition, the resulting Eigenvector corresponding to the smallest Eigenvalue yields a least square estimate $\mathbf{\widehat{\overline{X}}}$ of $\mathbf{\overline{X}}$ up to an unknown scale, which is recovered as the final row of $\mathbf{\widehat{\overline{X}}}$. Eq. 2 is equivalent to estimating the point $\widehat{\overline{p}}$ of $N_{k}\geq 4$ intersecting spheres centered respectively at $\overline{k}_{j}$ with radii $\hat{r}_{j}$, as in Fig. 1(a). A Hough solution to this sphere intersection problem was first proposed in [Wu], albeit in the discrete domain and therefore less accurate. Eq. 2 can be solved for all points $p$ that lie within an object’s visibility mask output from segmentation. This results in a 3D estimate of the visible surface of the object, expressed within the object frame, as shown in Fig. 1(b).
3.2 Network Architecture
The network architecture is illustrated in Fig. 2. The input to the model at training consists of the segmented object in the RGBD image $\mathbf{I}_{S}^{RGBD}$ , $N_{k}\geq 4$ ground truth keypoint coordinates $k_{\text{j}}$ and their corresponding segmented radial maps $\mathbf{R}_{S}$, and the ground truth normalized point cloud image $\mathbf{C}$ for the object in its current pose. For each object, the keypoints are either generated by KeyGNet [wu2024keygnet], or the proposed symmetric keypoints, as described in Sec. 3.3. The ground truth radial map data $\mathbf{R}_{S}$ is a $H\!\!\times\!\!W\!\!\times\!\!N_{k}$ tensor comprising the radial distances between the 3D scene point corresponding to each visible 2D image pixel, relative to the $N_{k}$ predefined 3D keypoints, where the 2D slice $\mathbf{R}_{S}^{j}=\mathbf{R}_{S}[:,:,j]$ represents the $j^{th}$ keypoint. The network’s output consists of the unsegmented estimate $\mathbf{\hat{R}}$ of $\mathbf{R}_{S}$.
The normalized object point cloud $\mathbf{C}$ is defined for an object’s canonical pose, with its values scaled to the range $[0,1]$, following the approach of NOCS [wang2019normalized] and Pix2Pose [Park]. When the model regresses $\mathbf{\hat{C}}$ to approximate $\mathbf{C}$, it provides an estimate of the object’s shape in the canonical space. Although not perfectly accurate due to regression errors and projection ambiguities, it preserves essential geometric structures. This approximation aids in interpreting object geometry for a given viewpoint pose and contributes to learning more accurate radial maps.
The network structure employs a ResNet-152 backbone, similar to PVNet [Peng], with two key differences. First, we replaced LeakyReLU with ReLU as the activation function because our radial voting scheme only includes positive values, unlike the vector voting scheme of PVNet which also required accommodating negative values. Second, we increased the number of skip connections linking the downsampling and upsampling layers from three to five, allowing for the inclusion of a richer collection of additional local features during upsampling [Long].
The loss includes radial map regression term $\mathcal{L}_{R}$, which is the mean absolute error of the estimated radial values:
$\mathcal{L}_{R}=\frac{1}{N}\sum_{p_{i}\in\mathbf{S}}\left(\left|\hat{r}_{i}-r_{i}\right|\right)$ (3)
with the summation taking place for radial values $r_{i}$ corresponding to all $N$ points $p_{i}$ in segmentation mask $\mathbf{S}$. There is also a soft $L1$ regression loss term $\mathcal{L}_{C}$ for the normalized object point cloud in the current pose, as proposed in [wang2019normalized]:
$\mathcal{L}_{C}=\frac{1}{N}\sum_{p_{i}\in\mathbf{S}}\left\{\begin{array}[]{ll}5\cdot(c_{i}-\hat{c_{i}})^{2}&\text{if }|c_{i}-\hat{c_{i}}|\leq 0.1\\ |c_{i}-\hat{c_{i}}|-0.05&\text{if }|c_{i}-\hat{c_{i}}|>0.1\end{array}\right.$ (4)
We also include a variation of the symmetric loss proposed in [wang2019normalized]. This loss minimizes the difference between the normalized coordinate map $\mathbf{C}$ with its estimated value $\hat{\mathbf{C}}_{S}$, for one pose in the pose symmetry set $\Theta$:
$\mathcal{L}_{P}=\min_{\Theta}(\frac{1}{N}\sum_{p_{i}\in\mathbf{S}}\left(\lfloor{\hat{c_{i}}\cdot n_{\text{b}}}\rfloor-\lfloor{c_{i}}\cdot n_{\text{b}}\rfloor\right)^{2})$ (5)
Unlike [wang2019normalized], we discretize the normalized coordinate values to fall within a coarse discrete range $[0,\ldots,n_{\text{b}}-1]$ for each of the three dimensions of normalized space. We found in practise that this discretization served to improve performance for nearly symmetric objects (such as the LINEMOD “glue” object), but not for truly symmetric objects (such as “eggbox”) which suffered from the regular ambiguities, as demonstrated in the ablation study in the Supplementary Material (Sec. S.5). For this reason, we call $\mathcal{L}_{P}$ the pseudo-symmetric loss. This loss term ensures that nearly symmetric objects are correctly aligned under the set $\Theta$ of pseudo-symmetric rotations, which improves the accuracy of radial regression.
The total loss $\mathcal{L}_{\text{total}}$ is calculated as a weighted sum of the three distinct components:
$\mathcal{L}_{\text{total}}=\lambda_{1}\cdot\mathcal{L}_{R}+\lambda_{2}\cdot\mathcal{L}_{C}+\lambda_{3}\cdot\mathcal{L}_{P}$ (6)
For all experiments, the scale values were set empirically to $\lambda_{1}\!=\!0.6$, $\lambda_{2}\!=\!0.2$ and $\lambda_{3}\!=\!0.2$ for asymmetric and pseudo-symmetric objects. For symmetric objects, $\lambda_{2}$ and $\lambda_{3}$ were set to zero. The relative impact of the three loss terms is shown in the Supplementary Material, (Sec. S.5).
During inference, the RGB image $\mathbf{I}^{RGB}$ is first processed by MaskRCNN [maskrcnn] to generate the segmented object mask $\hat{\mathbf{S}}$, as shown in Fig. 2. $\hat{\mathbf{S}}$ is then element-wise multiplied by $\mathbf{I}^{RGB}$ and depth image $\mathbf{I}^{D}$ to produce the segmented RGBD object image $\mathbf{I}_{\hat{S}}^{RGBD}$. This segmented image is fed into ResNet-152 to produce the radial map estimates $\hat{\mathbf{R}}$, and the segmented radial map estimate $\mathbf{\hat{R}}_{\hat{S}}$ is obtained through element-wise multiplication of $\hat{\mathbf{S}}$ and $\hat{\mathbf{R}}$.
Our proposed DLT Surface Estimation method then takes $\mathbf{\hat{R}}_{\hat{S}}$ along with known 3D keypoints in the object frame to estimate the object surface points, where for each pixel ${p}_{i}$ in $\mathbf{\hat{R}}_{\hat{S}}$, the estimated object surface point is denoted as $\hat{\bar{p}}_{i}$, and the full set of estimated surface points is represented as $\widehat{\overline{\mathbf{X}}}$. Each $\hat{\bar{p}}_{i}$ in the object frame has a corresponding $p_{i}$ in the image frame, both of which are indexed through the associated image pixel. To compute the full metric 6D pose of detected objects, we align $\mathbf{\widehat{\overline{X}}}$ with the input depth-derived 3D point cloud $\mathbf{I}_{\hat{S}}^{D}$ and estimate the 3D translation and rotation using a RANSAC-based Umeyama algorithm [umeyama, Fischler1981RandomSC] for an initial pose estimation, which may be further refined using ICP [icp].
3.3 Symmetric Keypoints
Earlier keypoint-based methods  [Wu, He, Peng] assign a fixed keypoint order without considering object symmetries, potentially leading to inconsistent keypoint assignments when multiple valid symmetric configurations exist. This ambiguity is particularly problematic for objects with discrete symmetries, such as those exhibiting $\pi$-radian symmetry, where the keypoints extracted in the image frame can map to a variety of respective corresponding keypoints in the object frame, depending on which equivalent symmetric pose is selected. Consequently, these methods struggle to consistently predict keypoints in a manner that preserves object symmetry, introducing regression ambiguities and errors which challenges the model to learn a stable representation.
Existing keypoint selection strategies, including bounding box keypoints [Wu], farthest point sampling keypoints [He], and KeyGNet keypoints [wu2024keygnet], attempt to improve the accuracy and robustness of the overall keypoint matching process. Among these, KeyGNet mitigates symmetry-related ambiguities by learning optimal keypoint locations rather than relying on heuristic selection, thereby enhancing stability. However, even with these improved keypoint locations, inconsistencies in ordering persist, as the keypoints still require correct indexing relative to the camera viewpoint for stable regression.
To address this challenge, we propose a symmetry-aware approach that enforces consistent keypoint ordering during training by dynamically adapting to object symmetries. The model estimates a four-channel radial map, where each channel encodes the radial distance of each image pixel to a predefined keypoint. Crucially, the structure of the radial map must remain consistent across symmetric object poses; otherwise, variations in keypoint ordering can lead to inconsistent regression targets and hinder generalization.
Unlike previous methods that rely on a fixed keypoint assignment, our approach dynamically reorders the radial channels based on the object’s symmetry and its orientation relative to the camera. Specifically, we determine the ordering of keypoints by analyzing their proximity to the camera origin. The approach depends on designing pairs of symmetric keypoints which correspond to the symmetric object poses. These keypoints are generated by leveraging the Oriented Bounding Box (OBB), which provides a stable reference frame for structured keypoint placement. The OBB is computed using the Minimum Volume Enclosing Box algorithm [CHAN20011433], ensuring a compact representation of the object. From the OBB, four side faces are identified, each defined by four corner points, and the face centers are computed as their mean. The face normal vectors are determined using the cross-product of two independent edge vectors and subsequently normalized. The symmetric keypoints are then generated by translating these face centers along their normal vectors by a fixed offset distance $d$. This ensures that all keypoints remain equidistant from the OBB center, preserving spatial symmetry. A pseudocode representation of the above described algorithm for generating symmetric keypoints is presented in Algorithm 1. To our knowledge, this the the first time that keypoint selection has been considered as an approach to address object symmetries.
Figure 3: Eggbox keypoints under different rotations. (a) Original pose. (b) 180-degree Z-axis rotation, demonstrating symmetry. (c) Top-down view, highlighting symmetric keypoints $k_{1}$-$k_{2}$ (top) and $k_{3}$-$k_{4}$ (bottom). (d) -45-degree Z-axis rotation. (e) 135-degree Z-axis rotation, equivalent to a 180-degree rotation of (d). (f) Side view, showing the spatial distribution of keypoints $k_{1}$ to $k_{4}$.
An example is shown in Fig. 3, in which the object exhibits a $\pi$-radian symmetry between pairs of keypoints $k_{1}$$k_{2}$ and $k_{3}$$k_{4}$, yielding a total of 4 possible orderings. For the configuration in Fig. 3(a), the four-channel radial map ($\mathbf{R}_{S}$ of Fig. 2) is ordered as $[\mathbf{R}_{S}^{2},\mathbf{R}_{S}^{1},\mathbf{R}_{S}^{4},\mathbf{R}_{S}^{3}]$ with the order determined by the keypoints’ relative proximity to the camera, i.e. $k_{2}$ is closer to the camera origin than $k_{1}$, and $k_{4}$ is closer than $k_{3}$. For Fig. 3(b), the correct ordering is $[\mathbf{R}_{S}^{1},\mathbf{R}_{S}^{2},\mathbf{R}_{S}^{3},\mathbf{R}_{S}^{4}]$ whereas in Fig. 3(d), the order is $[\mathbf{R}_{S}^{1},\mathbf{R}_{S}^{2},\mathbf{R}_{S}^{4},\mathbf{R}_{S}^{3}]$ while Fig. 3(e) follows order $[\mathbf{R}_{S}^{2},\mathbf{R}_{S}^{1},\mathbf{R}_{S}^{3},\mathbf{R}_{S}^{4}]$. Ensuring a consistent radial map order relative to the camera viewpoint prevents inconsistencies in loss computation, stabilizing training and improving pose estimation under symmetric transformations, as demonstrated in the ablation experiments of Sec. 4.5.
4 Evaluation
4.1 Datasets and Evaluation Metrics
Our method was trained and evaluated on three widely used public datasets, comprising LINEMOD (LM) [hinterstoisser2012model], LINEMOD Occlusion (LM-O) [Brachmann], and YCB-Video (YCB-V) [xiang2018posecnn]. The LM dataset consists of 15 sequences, each containing a single object with ground truth pose annotations. Each sequence includes approximately 1,200 images, totaling 18K images across the dataset. Following the standard protocol [xiang2018posecnn, Peng, bukschat2020efficientpose, kehl, wang2019normalized], we use 15% of the dataset for training and the remaining 85% for testing.
The LM-O dataset consists of 1,214 test images featuring 8 objects in partially occluded conditions, presenting a more challenging scenario. We use LM-O exclusively for evaluation and do not include it in training. Instead, we train our models on LM, utilizing approximately 1K real images per object. To further expand the training data for LM and LM-O, we incorporate physics-based rendering (PBR) data, which generates fully synthetic training images, and PVNet-rendering [Peng] augmentation, which further enhances the limited real images by overlaying objects from real images onto synthetic backgrounds.
YCB-V is a large-scale dataset consisting of 130K key frames captured across 92 video sequences featuring 21 objects. Following the training protocol of PVN3D [He], we use 113K frames for training and 27K frames for testing. Additionally, YCB-V includes 80K synthetic images where objects are rendered with randomized poses against a black background, which we incorporate for data augmentation.
Our primary evaluation metric is Average Recall ($AR$[hodan2018bop], which evaluates pose estimation performance across three key components: Visible Surface Discrepancy ($AR_{VSD}$), Maximum Symmetry-Aware Surface Distance ($AR_{MSSD}$), and Maximum Symmetry-Aware Projection Distance ($AR_{MSPD}$). These components enable a fine-grained analysis of pose accuracy while effectively handling object symmetries. For completeness, we also report results using the widely adopted ADD(-S) [hinterstoisser2012model] metric for the LM and LM-O datasets, as well as the ADD-S and ADD(-S) AUC [xiang2018posecnn] metric for the YCB-V dataset.
Method $\textit{AR}_{VSD}$ $\textit{AR}_{MSSD}$ $\textit{AR}_{MSPD}$ Mean AR
LM LM-O YCB-V LM LM-O YCB-V LM LM-O YCB-V LM LM-O YCB-V
PVNet [Peng] 0.502 0.683 0.730 0.638
ZebraPose [su2022zebrapose] 0.598 0.831 0.800 0.903 0.859 0.864 0.752 0.866
CosyPose [labbe2020] 0.670 0.567 0.831 0.810 0.748 0.903 0.849 0.826 0.850 0.773 0.714 0.861
RCVPose [Wu] 0.740 0.682 0.857 0.826 0.773 0.861 0.832 0.792 0.859 0.799 0.749 0.859
Pix2Pose [Park] 0.473 0.766 0.631 0.817 0.659 0.758 0.588 0.780
SurfEmb [Haugaard] 0.615 0.757 0.809 0.849 0.856 0.792 0.760 0.799
PFA [hu2022pfa] 0.658 0.863 0.843 0.920 0.890 0.881 0.797 0.888
CIR [lipson2022coupled] 0.601 0.871 0.778 0.924 0.824 0.885 0.734 0.893
DLTPose (Ours w/o ICP) 0.766 0.578 0.723 0.848 0.731 0.835 0.862 0.791 0.728 0.825 0.700 0.762
DLTPose (Ours) 0.801 0.678 0.879 0.891 0.834 0.922 0.903 0.877 0.884 0.865 0.797 0.895
Table 1: Performance comparison of leading 6DoF PE methods using Average Recall metrics on the LM, LM-O, and YCB-V datasets.
4.2 Implementation
Prior to training, RGB images are normalized to the range [0,1]. Segmented depth maps are independently normalized using their local minima and maxima to ensure consistent depth scaling across different scenes. Instead of computing radial distances directly from the depth field, radial maps are derived from the transformed object mesh and keypoints, ensuring consistency in distance estimation. These radial maps are expressed in decimeter units, as all LM, LM-O and YCB-V objects have a maximum diameter of 1.5 decimeters, allowing the network to operate within a stable numerical range and improving prediction accuracy [Wu].
For optimization, we use the Adam optimizer [adam] with an initial learning rate of 1e-3, following the loss functions in Eqs. (3)–(5). The learning rate is dynamically adjusted using a Reduce-on-Plateau strategy, where it is reduced by a factor of 0.1 upon stagnation. The method is implemented in PyTorch [pytorch], with each object being trained on a separate model for 200-250 epochs, using a batch size of 32.
Additionally, we train a custom Mask R-CNN [maskrcnn] model, pretrained on ImageNet [imgnet], to detect and segment the target objects in the scene. During training, RGB images are shifted and scaled to match the ImageNet mean and standard deviation. Separate models are trained for the LM, LM-O, and YCB-V datasets, each for $\sim\!\!50$ epochs with a batch size of 4 and an initial learning rate of 1e-4. All training is conducted on a server equipped with an Intel Xeon 5218 CPU and three RTX8000 GPUs or three A100 GPUs.
4.3 Results
Table 1 compares DLTPose with recent state-of-the-art methods on LM, LM-O, and YCB-V using Average Recall ($AR$). All methods refine initial estimated poses with a post processing step, typically ICP. Our method achieves the highest mean $AR$ across all datasets, with scores of 0.865 on LM, 0.797 on LM-O, and 0.895 on YCB-V.
For LM, DLTPose achieves 0.801 in $AR_{VSD}$, surpassing RCVPose (0.740) by +8.2% and CosyPose (0.670) by +19.5%. Additionally, our method obtains 0.891 in $AR_{MSSD}$, outperforming RCVPose (0.826) by +7.8% and CosyPose (0.810) by +10.0%. DLTPose also leads in $AR_{MSPD}$ with 0.903, improving over CosyPose (0.849) by +6.4% and RCVPose (0.832) by +8.5%. Consequently, our mean $AR$ score (0.865) outperforms RCVPose (0.799) by +8.3% and CosyPose (0.773) by +11.9%, establishing DLTPose as the current top-performing method on LM.
For LM-O, DLTPose achieves the highest Mean AR (0.797), outperforming PFA (0.797) and CIR (0.734) by +8.6%, demonstrating superior overall stability in occlusion-heavy settings. It also ranks second across $AR_{VSD}$ (0.678), $AR_{MSSD}$ (0.834), and $AR_{MSPD}$ (0.877). For YCB-V, DLTPose achieves the highest overall mean AR (0.895), surpassing CIR (0.893) by +0.2% and PFA (0.888) by +0.8%. While CIR leads in $AR_{MSPD}$ (0.885), slightly ahead of DLTPose (0.884) by -0.1%, our method ranks first in $AR_{VSD}$ (0.879), outperforming RCVPose (0.857) by +2.6%, and first in $AR_{MSSD}$ (0.922), surpassing CIR (0.924) by -0.2%.
Method LM LM-O YCB-V
-S (-S)
DenseFusion [wang2019densefusion] 94.3
REDE [Hua2020REDEEO] 98.9 65.4
MaskedFusion [pereira2020maskedfusion] 97.3
PR-GCN [Zhou2021PRGCNAD] 99.6 64.8
EANet [ea-net] 97.6 94.2
PVN3D [He] 99.4 63.3 96.1 92.3
RCVPose [Wu] 99.7 71.1 97.2
IRPE [Jin2024IRPEIR] 98.8 85.4 94.9 90.6
PoseCNN [xiang2018posecnn] 78.0 93.0 79.3
ZebraPose [su2022zebrapose] 78.5 92.0 87.5
DLTPose (Ours, w/o ICP) 99.0 80.3 98.3 93.9
DLTPose (Ours) 99.9 90.4 99.7 97.1
Table 2: Performance comparison of leading 6DoF PE methods using the ADD-S and ADD(-S) metrics, on the LM, LM-O, and YCB-V datasets.
In Table 2, we provide the mean ADD(-S) for LM and LM-O across all objects in the dataset for each approach, as well as the AUC for ADD-S and ADD(-S) on YCB-V. Detailed per-object results are provided in the Supplementary Material (Tables  S.3 – S.5). These results show that DLTPose consistently delivers top-tier performance across key metrics while maintaining stability across object categories.
4.4 Surface Estimation Accuracy Experiments
We conducted two experiments that show the improved accuracies and densities of surface points estimated with our DLT method, and its benefit on pose estimation. The first experiment compares the accuracies and number of surface points generated by leading dense approach SurfEmb [Haugaard] with our DLT approach, comparing them against ground truth values. Ground truth 2D-3D correspondences are obtained from the 3D mesh model, foreground mask, and object pose for each object in the LM-O dataset. From the known object pose, the 3D mesh is transformed into the camera frame. Each foreground point in the transformed mesh is then projected onto the image plane using the camera intrinsics, producing a ground truth set in which each 2D image pixel indexes a corresponding 3D object surface point.
Figure 4: Mean error vs. mean number of estimated surface points per object, for SurfEmb (red) and DLTPose (green). Solid lines average over all estimated points; dashed lines only include points with errors $\leq 10$ mm.
To assess accuracy of the estimates, we generated corresponding sets for both SurfEmb and our DLT approach and computed the point estimation error by comparing them against the ground truth correspondences. The errors between ground truth and estimated points for each method was measured across different percentile levels (i.e. top $10^{th},25^{th},50^{th},75^{th},$ and $100^{th}$ percentile) of the estimated correspondences, sorted by increasing error. The results are shown in Fig. 4, each curve plotting values from the smallest ($10^{th}$, leftmost) to the largest ($100^{th}$, rightmost) percentile. The solid curves include all estimated points, whereas the dashed curves include only points that lie within 10 mm of their ground truth corresponding values. These results demonstrate that the DLT approach outperforms SurfEmb in both accuracy and density of surface point estimates. As the number of estimated surface points increases, SurfEmb’s error rises sharpely, indicating higher deviations from the ground truth, whereas the DLT approach maintains lower errors across all cases. When filtering correspondences, and considering errors $\leq 10$ mm, both methods show improved accuracy, with the DLT approach retaining a greater number of higher accuracy surface points. Overall, the DLT approach achieves both increased accuracy and increased point count.
Figure 5: Average recall components (MSSD, MSPD, and $AR$) vs. standard deviation of injected noise on the estimated LM-O object surface points, both with (solid lines) and without (dotted lines) ICP refinement.
A second experiment showed the impact of higher accuracy surface estimates on pose estimation. Zero-mean Gaussian noise with varying standard deviations was added to the estimated object surface points, and the $AR$ metric was evaluated for all objects in the LM-O dataset following pose estimation, with the frontend RANSAC-enabled Umeyama used during inference. The results are plotted in Figure 5, and show that as noise increases, performance degrades across evaluation metrics MSSD, MSPD, and $AR$, reinforcing the impact of accurate surface point estimation on the accuracy of pose estimation.
4.5 Ablations
We performed two experiments, to identify the effect of symmetric keypoints, and the effect of the pseudo-symmetric loss on training. The details are described in the Supplementary Material, Sec. S.5. In summary, they show that the use of the symmetric keypoints (Sec. 3.3) improves both ADD-S and $AR$ scores for symmetric objects over KeyGNet keypoints (Table S.1), and that the pseudo-symmetric loss terms improve the ADD(-S) scores for all LM-O objects (Table S.2).
5 Conclusion
We have presented DLTPose, which estimates 3D surface points from a minimal set of four keypoints using a novel DLT formulation. The surface points are shown to be highly accurate, which leads to improved pose estimation accuracy. In addition, we present a symmetric keypoint ordering method that dynamically orders keypoints, thereby reducing ambiguities of regressed values during training.
Our method achieves state-of-the-art performance on the benchmark LM, LM-O and YCB-V datasets, outperforming recent methods, including leading dense methods such as SurfEmb. Its performance is particularly strong on symmetric objects, where the symmetric keypoint ordering approach improves accuracy.
Future work will focus on methods to further improve the accuracy of surface estimation, which will in turn improve pose estimation accuracy. We will also explore other approaches for generating and dynamically ordering symmetric keypoints during training.
Supplementary Material
S.1 Overview
This document provides additional details and experiments supporting our main work. Section S.2 presents the full derivation of the Direct Linear Transform (DLT) formulation. Section S.3 details our symmetric keypoint generation approach using Oriented Bounding Boxes (OBB). Section S.4 evaluates surface estimation accuracy by comparing error distributions for common pixel-aligned correspondences between our method and SurfEmb. Section S.5 covers ablation studies on symmetric keypoints, pseudo-symmetric loss, and surface estimation noise. Section S.6 provides per-object accuracy results for LINEMOD, Occlusion LINEMOD, and YCB-Video, including ADD(-S) and AUC scores, along with qualitative pose recovery examples.
S.2 DLT Derivation
Let $\bar{p}=(\bar{x},\bar{y},\bar{z})$ be an object surface point and let $\bar{k}=(\bar{x}_{k},\bar{y}_{k},\bar{z}_{k})$ be a keypoint, with $\bar{p}$ and $\bar{k}$ both described in a common reference frame, such as the object frame without loss of generality. Expanding Eq. 1, and for simplicity dropping the subscripts on $\bar{k}_{j}$ and $r_{j}$, the square of the radial distance $r$ between $\bar{p}$ and $\bar{k}$ is expressed as:
$(\bar{x}-\bar{x}_{k})^{2}+(\bar{y}-\bar{y}_{k})^{2}+(\bar{z}-\bar{z}_{k})^{2}=r^{2}.$ (S.1)
Expanding the terms of Eq. (S.1) gives:
$\bar{x}^{2}-2\bar{x}\bar{x}_{k}+\bar{x}_{k}^{2}+\bar{y}^{2}-2\bar{y}\bar{y}_{k}+\bar{y}_{k}^{2}+\bar{z}^{2}-2\bar{z}\bar{z}_{k}+\bar{z}_{k}^{2}-r^{2}=0,$ (S.2)
and further rearranging terms yields:
$-2\bar{x}\bar{x}_{k}-2\bar{y}\bar{y}_{k}-2\bar{z}\bar{z}_{k}+(\bar{x}^{2}+\bar{y}^{2}+\bar{z}^{2})+(\bar{x}_{k}^{2}+\bar{y}_{k}^{2}+\bar{z}_{k}^{2}-r^{2})=0.$ (S.3)
All known (i.e. constant or measured) quantities can be collecting into a left vector $A$, and multiplied with a right vector $X$ containing all unknown quantities, as follows:
$\begin{bmatrix}-2\bar{x}_{k}&-2\bar{y}_{k}&-2\bar{z}_{k}&1&\left(\bar{x}_{k}^{2}+\bar{y}_{k}^{2}+\bar{z}_{k}^{2}-r^{2}\right)\end{bmatrix}$ (S.4)
$\cdot\begin{bmatrix}\bar{x}\\ \bar{y}\\ \bar{z}\\ \left(\bar{x}^{2}+\bar{y}^{2}+\bar{z}^{2}\right)\\ 1\end{bmatrix}=0$
Finally, we simplify the notation to $|\!|\bar{k}|\!|^{2}=\bar{x}_{k}^{2}+\bar{y}_{k}^{2}+\bar{z}_{k}^{2}$ and $|\!|\bar{p}|\!|^{2}=\bar{x}^{2}+\bar{y}^{2}+\bar{z}^{2}$ and stack a series of $N_{k}$ such rows into matrix $A$ to yield Eq. 2:
$\begin{bmatrix}\!\!-2\bar{x}_{k_{1}}\!\!\!\!\!\!\!\!&\!\!\!\!-2\bar{y}_{k_{1}}\!\!\!\!\!\!\!\!&\!\!\!\!\!\!\!\!-2\bar{z}_{k_{1}}\!\!\!\!\!\!\!\!&\!\!\!\!1&\!\!\!\!(|\!|\bar{k}_{1}|\!|^{2}\!-\!\hat{r}_{1}^{2})\\ \vdots&\vdots&\vdots&\!\!\!\!\vdots&\vdots\\ \!-2\bar{x}_{k_{N_{k}}}\!\!\!\!\!\!\!\!&-2\bar{y}_{k_{N_{k}}}\!\!\!\!\!\!\!\!&-2\bar{z}_{k_{N_{k}}}\!\!\!\!&\!\!\!\!1&\!\!\!\!(|\!|\bar{k}_{N_{k}}|\!|^{2}\!-\!\hat{r}_{N_{k}}^{2})\\ \end{bmatrix}\begin{bmatrix}\bar{x}\\ \bar{y}\\ \bar{z}\\ |\!|\bar{p}|\!|^{2}\\ 1\end{bmatrix}=0.$
S.3 Symmetric Keypoints Generation
This method utilizes the Oriented Bounding Box (OBB), derived from the Minimum Volume Enclosing Box algorithm [CHAN20011433], to establish a consistent reference frame for keypoint placement. By identifying four primary side faces and computing their centers, the approach ensures a structured and uniform distribution of keypoints. The normal vectors of these faces, obtained through the cross-product of independent edge vectors, guide the placement process. Each center is shifted along its normal by a fixed offset ,̣ preserving equal spacing relative to the OBB center. This technique enforces stable keypoint assignments across symmetric object poses and the accompanying pseudocode 1 provides a clear implementation framework for integrating this approach into practical applications.
Algorithm 1 Symmetric Keypoint Generation from Oriented Bounding Box (OBB)
1:Object mesh $\mathcal{M}$, Offset distance $d$
2:Symmetric keypoints $\mathcal{K}=\{k_{1},k_{2},k_{3},k_{4}\}$
3:Step 1: Extract Oriented Bounding Box (OBB)
4:Compute OBB from $\mathcal{M}$ using Minimum Volume Enclosing Box.
5:Retrieve 8 corner points $\mathcal{Q}=\{q_{0},q_{1},\dots,q_{7}\}$.
6:Step 2: Define Side Faces
7:$F_{1}=\{q_{0},q_{1},q_{5},q_{4}\}=\{q_{0}^{1},q_{1}^{1},q_{2}^{1},q_{3}^{1}\}$
8:$F_{2}=\{q_{2},q_{3},q_{7},q_{6}\}=\{q_{0}^{2},q_{1}^{2},q_{2}^{2},q_{3}^{2}\}$
9:$F_{3}=\{q_{0},q_{2},q_{6},q_{4}\}=\{q_{0}^{3},q_{1}^{3},q_{2}^{3},q_{3}^{3}\}$
10:$F_{4}=\{q_{1},q_{3},q_{7},q_{5}\}=\{q_{0}^{4},q_{1}^{4},q_{2}^{4},q_{3}^{4}\}$
11:$\cal{K}=\{\}$
12:for all $F_{i}\in\{F_{1},F_{2},F_{3},F_{4}\}$ do
13:       Step 3: Compute Face Center
14:       Compute face center: $f_{i}=\frac{1}{4}\sum{\{q_{j}^{i}\}_{j=0}^{3}}$
15:       Step 4: Compute Face Normal Vector
16:       Compute edge vectors: $\vec{e}_{1}\!=\!q_{1}^{i}-q_{0}^{i},\;\vec{e}_{2}\!=\!q_{3}^{i}-q_{0}^{i}$
17:       Compute normal vector: $\vec{n}_{i}=\frac{\vec{e}_{1}\times\vec{e}_{2}}{||\vec{e}_{1}\times\vec{e}_{2}||}$
18:       Step 5: Compute Symmetric Keypoints
19:       Compute symmetric keypoint: $k_{i}=f_{i}+d\cdot\vec{n}_{i}$
20:       $\mathcal{K}=\mathcal{K}+k_{i}$
21:end for
S.4 Surface Estimation, Common Points
This experiment evaluates the accuracy of surface point estimates generated by SurfEmb [Haugaard] and our DLT approach, focusing specifically on those common pixels in the scene where both methods provided corresponding 3D surface estimates. The objective is to compare the estimation error for overlapping correspondences, ensuring a fair assessment of each method’s accuracy in shared regions.
The mean error per object in LMO is computed for these common pixels, and the results are visualized in Fig. S.1. The comparison shows that the DLT approach consistently achieves lower estimation error than SurfEmb across all density levels, demonstrating superior surface point accuracy even in overlapping regions. These findings reinforce that the DLT approach offers improved geometric consistency and more reliable surface reconstruction, ultimately leading to better pose estimation.
Figure S.1: Mean error vs. mean number of estimated surface points per object, for SurfEmb (red) and DLTPose (green). A common set of points are evaluated for each method at each percentile level ($10^{th},25^{th},50^{th},75^{th},100^{th}$), sorted by increasing error.
S.5 Ablation Experiments
Effect of Symmetric Keypoints on Pose Estimation. This experiment evaluates the impact of the proposed symmetric keypoint framework described in Sec. 3.3 on 6DoF pose estimation for objects with rotational symmetries. We compare the pose estimation results on the symmetric eggbox (from LM-O) and bowl (from YCB-V) objects, using the ADD-S and AR metrics. For each object, the keypoints were selected using either KeyGNet [wu2024keygnet] or our symmetric keypoint framework.
Table S.1 shows that our symmetric keypoints consistently improve performance. On eggbox, ADD-S increases from 86.5% to 92.0% (w/o ICP) and 97.8% to 98.6% (w/ ICP), while AR improves from 0.544 to 0.691. Similarly, bowl sees an increase in ADD-S from 88.8% to 99.56% (w/o ICP) and 97.9% to 100.0% (w/ ICP), with AR rising from 0.715 to 0.893. The YCB-V wood block shows further improvement, with ADD-S increasing from 99.5% to 100.0% (w/o ICP) and maintaining 100.0% (w/ ICP), while AR improves from 0.964 to 0.988 when using our symmetric keypoints. These results demonstrate that explicitly enforcing the proposed symmetric keypoint framework reduces regression ambiguities and improves pose estimation accuracy for symmetric objects. This is especially evident from the improvement in the $AR$ scores, which is more discriminating than the coarser ADD(-S) metric.
Object KeyGNet [wu2024keygnet] keypoints Symmetric keypoints
ADD-S ADD-S
w/o ICP w ICP $AR$ w/o ICP w ICP $AR$
LM-O eggbox 86.5 97.8 0.544 92.0 98.6 0.691
YCB-V bowl 88.8 97.9 0.715 99.5 100.0 0.893
YCB-V wood_block 99.5 100.0 0.964 100.0 100.0 0.988
Table S.1: Comparison of ADD-S (without and with ICP) and $AR$ metrics for two symmetric objects, using keypoints selected using either KeyGNet [wu2024keygnet] or our symmetric keypoint generation method (Sec. 3.3).
Effect of Pseudo-Symmetric Loss on Training. This ablation study examines the effect of different loss configurations on model performance by comparing training results using $\mathcal{L}_{R}$ (Eq. 3), a weighted combination of $\mathcal{L}_{R}$ and $\mathcal{L}_{C}$ ($\lambda_{1}\cdot\mathcal{L}_{R}+\lambda_{2}\cdot\mathcal{L}_{C}$, Eq. 4, with $\lambda_{1}=0.7,\lambda_{2}=0.3$), and the full loss $\mathcal{L}_{\text{total}}$ (Eq. 6). Table S.2 presents ADD(-S) results on LM-O across multiple objects, evaluating performance both with and without ICP refinement.
The results indicate that integrating pseudo-symmetric loss ($\mathcal{L}_{\text{total}}$) improves mean ADD(-S) by +3.6% without ICP and +1.9% with ICP. This demonstrates that enforcing pseudo-symmetry during training enhances radial map regression, leading to more accurate surface estimates and improved pose estimation accuracy.
Object w/o ICP w ICP
$\mathcal{L}_{R}$ $\mathcal{L}_{R}+\mathcal{L}_{C}$ $\cal{L}_{\text{total}}$ $\mathcal{L}_{R}$ $\mathcal{L}_{R}+\mathcal{L}_{C}$ $\cal{L}_{\text{total}}$
ape 73.5 73.3 74.2 85.5 86.5 87.1
can 88.2 89.1 93.3 94.2 96.2 96.8
cat 58.2 59.6 59.8 67.5 69.5 71.0
driller 87.6 88.1 90.6 95.5 95.1 97.3
duck 51.5 54.1 52.7 83.4 84.5 84.1
eggbox 83.4 86.4 86.5 96.9 97.6 97.8
glue 87.3 87.7 89.1 90.8 90.7 91.1
holepuncher 78.8 81.2 91.0 93.6 93.9 97.5
Mean 76.1 77.4 79.7 88.4 89.3 90.3
Table S.2: Comparison of ADD(-S) metric, without and with ICP, on LM-O for models trained with KeyGNet keypoints and three loss function configurations.
S.6 Accuracy Results Per Object
The detailed ADD(-S) results for the LM and LM-O datasets, along with the AUC results for ADD-S and ADD(-S) on the YCB-V dataset, are provided in Tables S.3S.5. Additionally, Figure S.2– S.4 presents qualitative examples of successful pose recoveries, where red dots represent projected surface points from ground truth poses, while blue dots correspond to those from the estimated poses.
As shown in Table S.3, the LM dataset is largely saturated, with multiple methods achieving near-perfect scores across most objects. However, our approach still achieves the highest mean ADD(-S) of 99.9%, with perfect scores (100%) on several objects, demonstrating its robustness in handling both symmetric and non-symmetric objects.
Table S.4 highlights the challenges posed by the LM-O dataset, where occlusions and imperfect object meshes introduce significant difficulties to pose estimation. Despite these challenges, our method achieves the highest mean accuracy of 90.4%, outperforming prior state-of-the-art approaches and showcasing strong robustness in occluded scenarios.
For the YCB-V dataset, Table S.5 reports AUC for both ADD-S and ADD(-S). Without ICP, our approach achieves a mean AUC of 98.3% for ADD-S and 93.9% for ADD(-S). With ICP, the performance further improves to 99.7% for ADD-S and 97.1% for ADD(-S), marking a significant improvement over prior state-of-the-art methods, i.e. +2.5% for ADD-S, and +4.8% for ADD(-S).
Figure S.2: Overlay results on select LM images.
Figure S.3: Overlay results on select LM-O images.
Figure S.4: Overlay results on select YCB-V images.
Object Dense- Fusion  [wang2019densefusion] RCVPose [Wu] PVN3D [He] REDE [Hua2020REDEEO] Masked- Fusion [pereira2020maskedfusion] PR-GCN [Zhou2021PRGCNAD] EANet [ea-net] IRPE [Jin2024IRPEIR] DLTPose w/o ICP DLTPose
ape 92.3 99.6 97.3 95.6 92.2 99.2 95.1 95.7 99.7 100.0
benchvise 93.2 99.7 99.7 99.4 98.4 99.8 97.5 99.9 99.8 100.0
camera 94.4 99.7 99.7 99.6 98.0 100.0 98.5 97.2 98.7 100.0
can 93.1 99.3 98.0 97.8 97.4 99.3 97.7 100.0 100.0 100.0
cat 96.5 99.7 99.9 99.5 97.7 99.8 97.7 99.6 99.8 100.0
driller 87.0 100.0 99.3 99.3 95.6 99.9 99.2 99.7 98.9 99.8
duck 92.3 99.7 99.4 98.0 94.0 98.2 97.3 97.9 100.0 100.0
eggbox 99.8 99.3 99.3 98.6 99.6 99.9 99.6 99.7 99.8 100.0
glue 100.0 99.7 100.0 100.0 100.0 100.0 99.9 99.7 100.0 100.0
holepuncher 92.1 100.0 99.3 98.6 97.3 99.0 96.8 98.3 99.8 99.9
iron 97.0 99.9 99.7 99.8 97.1 100.0 99.4 98.3 98.5 99.0
lamp 95.3 99.5 99.5 99.3 99.0 100.0 99.2 98.2 93.8 100.0
phone 92.8 99.7 99.5 99.5 98.8 99.9 98.7 98.2 97.8 100.0
Mean 94.3 99.7 99.4 98.9 97.3 99.6 97.6 98.8 99.0 99.9
Table S.3: Per object comparison of ADD(-S) results on LM dataset.
Object PoseCNN [xiang2018posecnn] PVN3D [He] RCVPose [Wu] REDE [Hua2020REDEEO] PR-GCN [Zhou2021PRGCNAD] ZebraPose [su2022zebrapose] IRPE [Jin2024IRPEIR] DLTPose w/o ICP DLTPose
ape 76.2 33.9 61.3 53.1 40.2 60.4 69.8 74.2 87.1
can 87.4 88.6 93.0 88.5 76.2 95.0 95.4 93.3 96.8
cat 52.2 39.1 51.2 35.9 57.0 62.1 72.4 59.8 71.0
driller 90.3 78.4 78.8 77.8 82.3 94.8 93.0 90.6 97.3
duck 77.7 41.9 53.4 46.2 30.0 64.5 81.1 52.7 84.1
eggbox* 72.2 80.9 82.3 71.8 68.2 73.8 85.3 92.0 98.6
glue* 76.7 68.1 72.9 75.0 67.0 88.7 88.4 89.1 91.1
holepuncher 91.4 74.7 75.8 75.5 97.2 88.4 97.9 91.0 97.5
Mean 78.0 63.2 71.1 65.4 64.8 78.5 85.4 80.3 90.4
Table S.4: Per object comparison of ADD or ADD-S results on LM-O. Asymmetric objects are evaluated with ADD, and symmetric objects (annotated with $\ast$) are evaluated with ADD-S.
Object PoseCNN [xiang2018posecnn] PVN3D [He] ZebraPose [su2022zebrapose] EANet  [ea-net] RCVPose [Wu] IRPE [Jin2024IRPEIR] DLTPose w/o ICP DLTPose
-S (-S) -S (-S) -S (-S) -S (-S) -S (-S) -S (-S) -S (-S) -S (-S)
002_master_chef_can 95.8 69.0 95.2 79.3 96.3 80.0 96.6 96.2 100.0 100.0 99.7 96.5 100.0 97.5
003_cracker_box 91.8 80.7 94.4 91.5 93.5 88.6 96.2 97.9 99.0 96.8 99.8 98.8 99.8 99.8
004_sugar_box 98.2 97.2 97.9 96.9 95.5 91.6 97.8 97.9 96.8 95.0 99.3 95.0 99.9 99.8
005_tomato_soup_can 94.5 94.3 95.9 89.0 94.4 90.3 96.1 99.0 98.2 94.1 100.0 99.6 100.0 99.8
006_mustard_bottle 98.4 87.0 98.3 97.9 96.1 93.0 96.9 98.2 99.0 95.3 100.0 99.9 100.0 100.0
007_tuna_fish_can 98.4 97.9 96.7 90.7 97.7 94.8 97.1 98.6 98.0 98.0 98.8 90.9 99.1 92.3
008_pudding_box 97.9 96.6 98.2 97.1 94.2 84.4 94.8 98.1 100.0 100.0 100.0 99.8 100.0 100.0
009_gelatin_box 98.8 96.6 98.8 98.3 97.8 88.4 97.1 98.4 99.0 98.9 99.8 97.0 100.0 100.0
010_potted_meat_can 92.8 83.8 93.8 87.9 93.8 84.0 97.2 98.4 86.0 74.4 99.1 97.4 99.7 98.1
011_banana 96.9 92.6 98.2 96.0 92.3 84.3 97.1 98.3 99.9 98.9 96.9 89.9 99.9 97.1
019_pitcher_base 97.8 92.3 97.6 96.9 89.8 89.0 98.0 97.2 91.2 87.8 98.3 84.7 99.5 87.9
021_bleach_cleanser 96.8 92.3 97.2 95.9 89.8 89.0 98.0 99.6 79.3 74.0 98.8 90.3 99.8 96.9
[…]
025_mug 95.1 91.1 97.7 96.0 99.9 99.9 97.6 98.7 99.7 99.7 98.6 84.3 99.4 88.3
035_power_drill 98.3 73.1 97.1 95.7 95.8 81.8 94.3 96.4 98.2 96.4 97.4 89.2 99.8 98.1
[…]
037_scissors 92.2 84.8 92.4 95.0 87.2 91.9 94.0 96.4 99.7 99.7 80.1 67.4 97.1 87.0
040_large_marker 97.2 47.3 98.1 91.6 97.6 89.7 94.0 96.6 97.6 89.7 99.9 95.2 99.9 96.1
[…]
Mean 93.0 79.3 96.1 92.3 92.0 87.5 94.2 97.2 94.9 90.6 98.3 93.9 99.7 97.1
Table S.5: Per object comparison of AUC of ADD-S/ADD(-S) results on YCB-V dataset.